An electron is moving through a magnetic field with a magnitude of 0.000045 T. If the electron is moving at a speed of 200,000 m/s, what is the magnetic force on the particle?
1
Expert's answer
2019-08-20T10:24:38-0400

"B =4.5{\cdot}10^{-5}T"

"v=2{\cdot}10^{5}{\frac{m}{s}}"

"q=1.6{\cdot}10^{-19}C"

Lorentz force:

"F=q{\cdot}v{\times}B=1.6{\cdot}10^{-19}{\cdot}2{\cdot}10^{5}{\times}4.5{\cdot}10^{-5}=
=14.4{\cdot}10^{-19}(N)"